Grade 11Algebra

how the rectangular hyperbola equation :xy=c^2 comes

Nusrat Arohi
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
We realize that the asymptotes of the hyperbola x2/a2 – y2/b2 = 1 … (1) 
are given by y = + (b/a) x … (2) 
On the off chance that θ be the edge between the asymptotes, then 
θ = tan–1 ((m1–m2)/(1 + m1m2)) 
= tan–1 [{(b/a)–(–b/a)}/{1+(b/a)(–b/a)}] 
= tan–1 [2(b/a)/(1–(b2/a2))] 
= 2 tan–1 (b/a) 
Be that as it may, if the hyperbola is rectangular, then θ = π/2 
i.e., π/2 = 2 tan–1 (b/an) or tan (π/4) = b/a 
⇒ b = a 
Consequently, from (1) the condition of the rectangular hyperbola is x2 – y2 = a2. 
With a specific end goal to get the condition of the hyperbola which has asymptotes as facilitate hub we pivot the tomahawks of reference through an edge of - 45o. 
Subsequently, for this we need to compose x/√2 + y/√2 for x and –x/√2 + y/√2 for y. 
The condition (i) progresses toward becoming 
(1/2)(x + y)2 – (1/2)(x – y)2 = a2 i.e. 
xy = ½ a2 or xy = c2 where c2 = a2/2.